OpenStudy (ktbug3):

Angie read 3 books in 5 days, and each book contained 405 pages. If Angie read the same number of pages each day, how many pages did she read per day? 89 1,215 243 405

OpenStudy (ktbug3):

@Camthecam @hakunamatata34 @skullpatrol @kxttxn

OpenStudy (phi):

you want \( \frac{pages}{day}\) you are given \( \frac{books}{day}\) Here is the idea: multiply by \(\frac{pages}{book}\) like this \[ \frac{books}{day} \cdot \frac{pages}{book} \] the "books" cancel, and we get \[ \frac{\cancel{books}}{day} \cdot \frac{pages}{\cancel{book}}= \frac{pages}{day}\]

OpenStudy (phi):

So, first step, using read 3 books in 5 days, can you write the fraction that stands for books per day ?

OpenStudy (ktbug3):

3/5?

OpenStudy (phi):

yes next, what is the fraction for pages per book ? use each book contained 405 pages.

OpenStudy (ktbug3):

405/3

OpenStudy (phi):

almost, it's 405/1 (that says 405 pages for 1 book)

OpenStudy (ktbug3):

o ok

OpenStudy (phi):

now you do the multiplication \[ \frac{3}{5} \cdot \frac{405}{1} \] (though the 1 in the 2nd fraction is not usually written)

OpenStudy (ktbug3):

1215/5

OpenStudy (phi):

I would first divide the 5 into 405 then multiply by 3

OpenStudy (ktbug3):

243

OpenStudy (phi):

and if you write that as a fraction it would be \[ \frac{243}{1} \] and if we keep track of the labels it would be \[ \frac{243 \ pages}{1 \ day} \]

OpenStudy (phi):

which can also be written as 243 pages per day

OpenStudy (ktbug3):

is that the answer?

OpenStudy (ktbug3):

thnx